OpenStudy (anonymous):

Find the value of c such that the point (2, 5) lies on the graph of f(x) = cx^2 + 4x − 2. c=

jimthompson5910 (jim_thompson5910):

All points are of the form (x,y) So, (x,y) = (2, 5) which means x = 2, y = 5

jimthompson5910 (jim_thompson5910):

So the idea is that if you plug in x = 2, you should get y = 5 or f(x) = 5

jimthompson5910 (jim_thompson5910):

f(x) = cx^2 + 4x − 2 f(x) = c(2)^2 + 4(2) − 2 ... plug in x = 2 5 = c(2)^2 + 4(2) − 2 ... plug in f(x) = 5 now solve for c

OpenStudy (anonymous):

so c=2?

jimthompson5910 (jim_thompson5910):

c = 2 is incorrect

OpenStudy (anonymous):

would the answer be -5?

jimthompson5910 (jim_thompson5910):

5 = c(2)^2 + 4(2) − 2 5 = c(4) + 4(2) − 2 5 = 4c + 8 - 2 5 = 4c + 6 c = ??

OpenStudy (anonymous):

-1/4?

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

Thanks! people are very helpful